OpenStudy (anonymous):

Derivative of t^2+2t-3

OpenStudy (lgbasallote):

2t + 2

myininaya (myininaya):

\[(t^n)'=nt^{n-1}\] \[(ct)'=c(t)'=c(1)=c\] where c is constant \[(c)'=0\]